OpenStudy (anonymous):

Determine the domain and range of the function: f(x)=4e^(x+2)+5

OpenStudy (anonymous):

\[e^x \\ Domain \ : \ \mathbb{R} \\Range \ : \ (0,\infty)\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so i change that answer based on the rest of the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so would the domain be x>5?

OpenStudy (anonymous):

or would it be x<-2?

OpenStudy (anonymous):

\[-\infty<x+2<\infty \ \ then \ \ -\infty<x<\infty\]

OpenStudy (anonymous):

the domain is R again

OpenStudy (anonymous):

Range is x>5